# Biased Backpressure Routing Using Link Features and Graph Neural Networks

### v1

@@ VII-B Network Size and Traffic Type
VII-B1 Streaming traffic
When the traffic is configured with all streaming flows, the average end-to-end delays (unit: time slots) as a function of the network size for ten tested BP schemes are presented in Fig. 5(a). Although not presented in Figs. 5, the ranking of the BP schemes on packet delivery rate is consistent with the inverted ranking on latency, showing that the latency under streaming traffic is dominated by the contribution of undelivered packets, whose delays are counted as $T$. This is consistent with the results under mixed traffic shown in Table II (Section VII-B3).
@@ VII-B2 Bursty traffic
Under lighter bursty traffic, both HOL and expQ can significantly improve the latency of EDR-$\bar{r}$ (Fig. 5(b)), boosting its delivery rate from $91.0\%$ to $99.8\%$ and $99.5\%$, respectively [2]. However, as the flow rates triple, they degrade the latency of EDR-$\bar{r}$ (Fig. 5(c)) but still improve its delivery rate (Fig. 5(d)). Such an increase in congestion is caused by the fact that HOL and expQ can reduce the network capacity of biased BP, as discussed in Section VII-C. A similar effect can be observed when expQ is applied to SP-$\bar{r}/(xr)$. Under lighter bursty traffic (Fig. 5(b)), SP-$\bar{r}/(xr)$-min, SP-$\bar{r}/(xr)$, and SP-$\bar{r}/(xr)$-expQ achieve the 1st, 2nd, and 3rd best latency, at a $100\%$ delivery rate. However, under heavier bursty traffic, as shown in Figs. 5(c) and 5(d), expQ worsens the latency and delivery rate of SP-$\bar{r}/(xr)$, while optimal bias scaling improves both performances. For biased BP, expQ consistently outperforms HOL and SJB, however, the effectiveness of these delay-based backlog metrics is limited to lighter traffic.
VII-B3 Mixed Traffic
We next evaluate low-overhead BP schemes under a mixture of streaming and bursty flows, to assess their effectiveness on the last packet problem under the interference of streaming traffic. In this case, a source-destination pair is configured as either streaming or bursty flow at equal probabilities. The bursty flows are configured with the heavier traffic loads described in Section VII-B2, to ensure similar numbers of packets are injected into the network for both streaming and bursty flows. We present the average latency and packet delivery rates of streaming and bursty flows under 15 tested BP schemes, alongside their rankings in Table II. In general, their rankings are consistent with the results in Sections VII-B1 and VII-B2 with similar traffic loads, with the exception that expQ is able to improve EDR-$\bar{r}$ and SP-$\bar{r}/(xr)$ for bursty flows in mixed traffic.
@@ VII-B3 Mixed Traffic
VII-C Network Capacity Region
To illustrate the impact of different low-overhead BP schemes on network capacity, we test them on 10 random networks of size $|{\mathcal{V}}|=100$ under streaming flows with constant arrival rates ranging from $0.1\leq\lambda\leq 12$. The total number of packets delivered to their destinations is collected for each test instance after $T=1000$ time steps, and the average end-to-end total throughput of 15 BP schemes as a function of flow arrival rate are presented in Fig. 7 and the corresponding peak values are listed in Table II (last column). For most BP schemes, the end-to-end total throughput first increases with the flow arrival rate, and then decreases slowly after reaching a peak. When the flow arrival rate is too large, the MaxWeight scheduler will prioritize the links near source nodes, so that fewer packets will reach their destinations, causing a decrease in the end-to-end throughput.
@@ VII-C Network Capacity Region
VII-D Node mobility
TABLE III: End-to-end delay on random networks of $100$ nodes, $T=1000$ under the unit-disk interference model and node mobility.
EDR-$\bar{r}$ SP-$\bar{r}/(xr)$
Bias update Delay (std.) Delivery (std.) Delay (std.) Delivery (std.)
@@ VII-D Node mobility
Lastly, we run 100 test instances on 10 draws of random networks with 100 nodes and node mobility, in which for every 100 time steps, $10$ random nodes take a random step on the 2D plane, i.e., $\bm{\kappa}\in{\mathbb{R}}^{2},\bm{\kappa}_{i}\sim\mathbb{N}(0,0.1)$ (normal distribution with mean of zero and standard deviation of $0.01$), conditioned on keeping the network connected. The long-term link rate of a newly created link $e$ is set as $r_{e}\sim\mathbb{U}(10,42)$. We compare the end-to-end delay and delivery rate of our neighborhood bias update rule in (10) with those of ideal (instantaneous) SSSP in Table III. Under the practical neighborhood update rule, SP-$\bar{r}/(xr)$ suffers a larger loss from the ideal (but impractical) SSSP than EDR-$\bar{r}$, i.e., the delivery rate of SP-$\bar{r}/(xr)$ dropped by $12\%$ compared to a drop of $5.6\%$ with EDR-$\bar{r}$. Despite larger delays and lower delivery rates, EDR-$\bar{r}$ suffers less under node mobility, showing that it is more robust to transient errors in biases.
VIII Conclusions
@@ Appendix A Proof of Inequality
Appendix B Justification of lightweight traffic setting
For all streaming flow setting, given the test setting in Section VII, in each test instance, the expected number of source nodes is $0.4|{\mathcal{V}}|$, each with an average arrival rate of $0.6=(0.2+1.0)/2$, the average exogenous packet arrival rate per node is computed as $0.24=0.6\times 0.4$.

### v2

@@ 7.2 NETWORK SIZE AND TRAFFIC TYPE
7.2.1 STREAMING TRAFFIC
When the traffic is configured with all streaming flows, the average end-to-end delays (unit: time slots) as a function of the network size for ten tested BP schemes are presented in Fig. 5(a). Although not presented in Figs. 5, the ranking of the BP schemes on packet delivery rate is consistent with the inverted ranking on latency, showing that the latency under streaming traffic is dominated by the contribution of undelivered packets, whose delays are counted as $T$. This is consistent with the results under mixed traffic shown in Table 2 (Section 7.2.3).
@@ 7.2.2 BURSTY TRAFFIC
Under lighter bursty traffic, both HOL and expQ can significantly improve the latency of EDR-$\bar{r}$ (Fig. 5(b)), boosting its delivery rate from $91.0\%$ to $99.8\%$ and $99.5\%$, respectively [2]. However, as the flow rates triple, they degrade the latency of EDR-$\bar{r}$ (Fig. 5(c)) but still improve its delivery rate (Fig. 5(d)). Such an increase in congestion is caused by the fact that HOL and expQ can reduce the network capacity of biased BP, as discussed in Section 7.3. A similar effect can be observed when expQ is applied to SP-$\bar{r}/(xr)$. Under lighter bursty traffic (Fig. 5(b)), SP-$\bar{r}/(xr)$-min, SP-$\bar{r}/(xr)$, and SP-$\bar{r}/(xr)$-expQ achieve the 1st, 2nd, and 3rd best latency, at a $100\%$ delivery rate. However, under heavier bursty traffic, as shown in Figs. 5(c) and 5(d), expQ worsens the latency and delivery rate of SP-$\bar{r}/(xr)$, while optimal bias scaling improves both performances. For biased BP, expQ consistently outperforms HOL and SJB, however, the effectiveness of these delay-based backlog metrics is limited to lighter traffic.
7.2.3 MIXED TRAFFIC
We next evaluate low-overhead BP schemes under a mixture of streaming and bursty flows, to assess their effectiveness on the last packet problem under the interference of streaming traffic. In this case, a source-destination pair is configured as either streaming or bursty flow at equal probabilities. The bursty flows are configured with the heavier traffic loads described in Section 7.2.2, to ensure similar numbers of packets are injected into the network for both streaming and bursty flows. We present the average latency and packet delivery rates of streaming and bursty flows under 15 tested BP schemes, alongside their rankings in Table 2. In general, their rankings are consistent with the results in Sections 7.2.1 and 7.2.2 with similar traffic loads, with the exception that expQ is able to improve EDR-$\bar{r}$ and SP-$\bar{r}/(xr)$ for bursty flows in mixed traffic.
@@ 7.2.3 MIXED TRAFFIC
7.3 NETWORK CAPACITY REGION
To illustrate the impact of different low-overhead BP schemes on network capacity, we test them on 10 random networks of size $|{\mathcal{V}}|=100$ under streaming flows with constant arrival rates ranging from $0.1\leq\lambda\leq 12$. The total number of packets delivered to their destinations is collected for each test instance after $T=1000$ time steps, and the average end-to-end total throughput of 15 BP schemes as a function of flow arrival rate are presented in Fig. 7 and the corresponding peak values are listed in Table 2 (last column). For most BP schemes, the end-to-end total throughput first increases with the flow arrival rate, and then decreases slowly after reaching a peak. When the flow arrival rate is too large, the MaxWeight scheduler will prioritize the links near source nodes, so that fewer packets will reach their destinations, causing a decrease in the end-to-end throughput.
@@ 7.3 NETWORK CAPACITY REGION
7.4 NODE MOBILITY
Table 3: End-to-end delay on random networks of $100$ nodes, $T=1000$ under the unit-disk interference model and node mobility.
Bias update EDR-$\bar{r}$ SP-$\bar{r}/(xr)$
Delay (std.) Delivery (std.) Delay (std.) Delivery (std.)
@@ 7.4 NODE MOBILITY
Lastly, we run 100 test instances on 10 draws of random networks with 100 nodes and node mobility, in which for every 100 time steps, $10$ random nodes take a random step on the 2D plane, i.e., $\bm{\kappa}\in{\mathbb{R}}^{2},\bm{\kappa}_{i}\sim\mathbb{N}(0,0.1)$ (normal distribution with mean of zero and standard deviation of $0.01$), conditioned on keeping the network connected. The long-term link rate of a newly created link $e$ is set as $r_{e}\sim\mathbb{U}(10,42)$. We compare the end-to-end delay and delivery rate of our neighborhood bias update rule in (10) with those of ideal (instantaneous) SSSP in Table 3. Under the practical neighborhood update rule, SP-$\bar{r}/(xr)$ suffers a larger loss from the ideal (but impractical) SSSP than EDR-$\bar{r}$, i.e., the delivery rate of SP-$\bar{r}/(xr)$ dropped by $12\%$ compared to a drop of $5.6\%$ with EDR-$\bar{r}$. Despite larger delays and lower delivery rates, EDR-$\bar{r}$ suffers less under node mobility, showing that it is more robust to transient errors in biases.
8 CONCLUSIONS
@@ 9 PROOF OF INEQUALITY
10 JUSTIFICATION OF LIGHTWEIGHT TRAFFIC SETTING
For all streaming flow setting, given the test setting in Section 7, in each test instance, the expected number of source nodes is $0.4|{\mathcal{V}}|$, each with an average arrival rate of $0.6=(0.2+1.0)/2$, the average exogenous packet arrival rate per node is computed as $0.24=0.6\times 0.4$.